OpenStudy (anonymous):

Solve a triangle: b = 6, c = 8, B = 75° One, two, or no triangles. Solve. My calculator is getting a Domain error when solving for C. Is there no triangle here?

OpenStudy (anonymous):

are you using Law of sine?

OpenStudy (anonymous):

Yes.  So, \[\sin 75\div6 = \sin c \div 8\] for angle C \[(8 \sin 75) \div 6\] However there is a domain error for sin here, obviously.

OpenStudy (anonymous):

Can noone help me?

OpenStudy (anonymous):